n Munich a bratwurst costs 6 euros; a hot dog costs $6.00 at Bostons Fenway Park. At an exchange rate of $1.60/per euro, what is the price of a bratwurst in terms of hot dogs? The relative price of bratwurst is hot dogs (Enter your response rounded to two decimal places) All else equal, how does this relative price change if the dollar appreciates to $1.25? Now the relative price of bratwurst ishot dogs (Enter your response rounded to two decimal places.) Compared with the initial situation, a hot dog has becomee expensive relative to a bratwurst
Answer #1

Price of 1 bratwurst = 6 euros x 1.60 = $9.6

Relative price of bratwurst = 9.6/6 =1.6 hot dogs

When the exchange rate is 1.25

Price of one bratwurst = 6x1.25 = $7.5

Relative price of bratwurst = 7.5/6 = 1.25 hot dogs

Compared with the initial situation,a hot dog has become more expensive relative to a bratwurst